# Testing the heavy decaying sterile neutrino hypothesis at the DUNE near detector

### v1

@@ I Introduction
The paper is organized as follows. In Section II, we introduce the heavy decaying sterile neutrino scenario considered in this paper and provide the main formulae used in our analysis. Section III briefly presents the specifications of the DUNE liquid argon near detector and describes the simulations performed in our paper. In Section IV, we present our results for the expected sensitivity of ND-LAr to the HDSN parameters, and comment about the possibility to distinguish between Dirac and Majorana neutrinos if this scenario is realized in Nature. Finally, we give our conclusions in Section V.
II The heavy decaying sterile neutrino (HDSN) scenario
@@
$P_{\mu\mu}=P_{\bar{\mu}\bar{\mu}}=(1-|U_{\mu 4}|^{2})^{2}+|U_{\mu 4}|^{4}\,e^{-\Gamma_{4}L}\,,\qquad P_{ee}=P_{\bar{e}\bar{e}}=1\ ,$ (6)
$P_{\mu e}=P_{\bar{\mu}\bar{e}}=|U_{\mu 4}|^{2}(1-e^{-\Gamma_{4}L})\ .$ (7)
@@
$\frac{dP_{\mu e}}{dE_{e}}=\frac{dP_{\bar{\mu}\bar{e}}}{dE_{e}}=|U_{\mu 4}|^{2}(1-e^{-\Gamma^{\rm D}_{4}L})\,\frac{2E_{e}}{E^{2}_{4}}$ ${\rm(Dirac\ case})\ ,$ (9)
$\frac{dP_{\mu e}}{dE_{e}}=\frac{dP_{\bar{\mu}\bar{e}}}{dE_{e}}=\frac{dP_{\mu\bar{e}}}{dE_{e}}=\frac{dP_{\bar{\mu}e}}{dE_{e}}=|U_{\mu 4}|^{2}(1-e^{-\Gamma^{\rm M}_{4}L})\,\frac{E_{e}}{E^{2}_{4}}$ ${\rm(Majorana\ case})\ .$ (10)
@@
III Experimental setup and Simulation details
@@ III.2 Simulation details
Table 1: Event rates at the DUNE liquid argon near detector (ND-LAr) expected in the Standard Model, assuming 3.5 years of on-axis data taking in either neutrino or antineutrino mode. The first and second rows correspond to the appearance channel in the neutrino and antineutrino modes, respectively, for which the signal from neutrino oscillations is negligible. The last three columns show the number of intrinsic, misidentification, and neutral current background events, as defined in Section II. The third and fourth rows correspond to the disappearance channel in the neutrino and antineutrino modes, respectively, and the number of signal events is given separately for the right sign component (e.g. $\nu_{\mu}\rightarrow\nu_{\mu}$ in the neutrino mode) and for the wrong sign component (e.g. $\bar{\nu}_{\mu}\rightarrow\bar{\nu}_{\mu}$ in the neutrino mode) of the neutrino beam.
@@
$\left[N^{\nu_{\mu}\rightarrow\nu_{e}}_{i}\right]^{D,\,CC}$ $=\frac{nT}{4\pi L^{2}}\,\int_{E_{rec}^{i}}^{E_{rec}^{i+1}}\!\!A(E_{rec})\,dE_{rec}\int_{0}^{E_{\nu}^{max}}\!\!\!dE_{\nu_{e}}\,\mathcal{R}(E_{\nu_{e}},\,E_{rec})\,\sigma^{CC}_{\nu_{e}}(E_{\nu_{e}})$
@@
$\left[N^{\nu_{\mu}\rightarrow\nu_{\mu}}_{i}\right]^{D,\,CC}$ $=\frac{nT}{4\pi L^{2}}\,\int_{E_{rec}^{i}}^{E_{rec}^{i+1}}\!\!A(E_{rec})\,dE_{rec}\int_{0}^{E_{\nu}^{max}}\!\!\!dE_{\nu_{\mu}}\,\mathcal{R}(E_{\nu_{\mu}},\,E_{rec})\,\sigma^{CC}_{\nu_{\mu}}(E_{\nu_{\mu}})\,\phi_{\nu_{\mu}}(E_{\nu_{\mu}})\,P^{D}_{\mu\mu}(E_{\nu_{\mu}})\,$, (19)
$\left[N^{\nu_{\mu}}_{i}\right]^{D,\,NC}$ $=\frac{nT}{4\pi L^{2}}\,\int_{E_{rec}^{i}}^{E_{rec}^{i+1}}\!\!A(E_{rec})\,dE_{rec}\ \biggl{[}\int_{0}^{E_{\nu}^{max}}\!\!\!dE_{\nu_{\mu}}\,\mathcal{R}(E_{\nu_{\mu}},E_{rec})\,\sigma^{NC}_{\nu}(E_{\nu_{\mu}})\,\phi_{\nu_{\mu}}(E_{\nu_{\mu}})\,P^{D}_{\mu\mu}(E_{\nu_{\mu}})$
$+\int_{0}^{E_{\nu}^{max}}\!\!\!dE_{\nu_{e}}\,\mathcal{R}(E_{\nu_{e}},E_{rec})\,\sigma^{NC}_{\nu}(E_{\nu_{e}})\int_{E_{\nu_{e}}}^{E_{\nu}^{max}}\!\!\!dE_{\nu_{\mu}}\,\phi_{\nu_{\mu}}(E_{\nu_{\mu}})\,\frac{dP^{D}_{\mu e}}{dE_{\nu_{e}}}\,(E_{\nu_{\mu}},E_{\nu_{e}})\biggr{)}\biggr{]}\ .$ (20)
@@
$\left[N^{\nu_{\mu}\rightarrow\nu_{\mu}}_{i}\right]^{M,\,CC}=\frac{nT}{4\pi L^{2}}\,\int_{E_{rec}^{i}}^{E_{rec}^{i+1}}\!\!A(E_{rec})\,dE_{rec}\int_{0}^{E_{\nu}^{max}}\!\!\!dE_{\nu_{\mu}}\,\mathcal{R}(E_{\nu_{\mu}},E_{rec})\,\sigma^{CC}_{\nu_{\mu}}(E_{\nu_{\mu}})\,\phi_{\nu_{\mu}}(E_{\nu_{\mu}})\,P^{M}_{\mu\mu}(E_{\nu_{\mu}})\,$, (22)
$\left[N^{\nu_{\mu}}_{i}\right]^{M,\,NC}$ $=\frac{nT}{4\pi L^{2}}\,\int_{E_{rec}^{i}}^{E_{rec}^{i+1}}\!\!A(E_{rec})\,dE_{rec}\ \biggl{[}\int_{0}^{E_{\nu}^{max}}\!\!\!dE_{\nu_{\mu}}\,\mathcal{R}(E_{\nu_{\mu}},E_{rec})\,\sigma^{NC}_{\nu}(E_{\nu_{\mu}})\,\phi_{\nu_{\mu}}(E_{\nu_{\mu}})\,P^{M}_{\mu\mu}(E_{\nu_{\mu}})$
@@
It is straightforward to generalize Eqs. (18) to (23) to the case where the experiment runs in the antineutrino mode. The probabilities appearing in these equations are given by (see Section II)
$\frac{dP^{D}_{\mu e}}{dE_{e}}=\frac{dP^{D}_{\bar{\mu}\bar{e}}}{dE_{e}}=|U_{\mu 4}|^{2}(1-e^{-\Gamma^{\rm D}_{4}L})\,\frac{2E_{e}}{E^{2}_{4}}$ ${\rm(Dirac\ case})\ ,$ (24)
$\frac{dP^{M}_{\mu e}}{dE_{e}}=\frac{dP^{M}_{\bar{\mu}\bar{e}}}{dE_{e}}=\frac{dP^{M}_{\mu\bar{e}}}{dE_{e}}=\frac{dP^{M}_{\bar{\mu}e}}{dE_{e}}=|U_{\mu 4}|^{2}(1-e^{-\Gamma^{\rm M}_{4}L})\,\frac{E_{e}}{E^{2}_{4}}$ ${\rm(Majorana\ case})\ ,$ (25)
@@
$P^{D}_{\mu\mu}=P^{D}_{\bar{\mu}\bar{\mu}}=\left(1-|U_{\mu 4}|^{2}\right)^{2}+|U_{\mu 4}|^{4}e^{-\Gamma^{\rm D}_{4}L}$ ${\rm(Dirac\ case})\ ,$ (26)
$P^{M}_{\mu\mu}=P^{M}_{\bar{\mu}\bar{\mu}}=\left(1-|U_{\mu 4}|^{2}\right)^{2}+|U_{\mu 4}|^{4}e^{-\Gamma^{\rm M}_{4}L}$ ${\rm(Majorana\ case})\ .$ (27)
IV ND-LAr sensitivity to the HDSN parameters
@@
IV.1 Appearance and disappearance spectra in the Standard Model
@@
IV.2 Appearance and disappearance spectra in the HDSN scenario
@@
IV.3 Sensitivity of ND-LAr to the HDSN parameters
@@
The different shapes of the appearance and disappearance sensitivity curves in the left plot of Fig. 5 can be understood from the formulae of Section II. For the disappearance channel, the difference between the predictions of the SM and HDSN scenario for the signal and background scales as $|U_{\mu 4}|^{2}$, as already noticed, with little dependence on $gm_{4}$ (see Eqs. (6) and (12), where the $gm_{4}$ dependence is hidden in the sterile neutrino decay rate $\Gamma_{4}$). In fact, the sensitivity to $|U_{\mu 4}|^{2}$ in the disappearance channel is almost independent of $gm_{4}$, except around $gm_{4}\sim 10\,\mbox{eV}$. The increased sensitivity to $|U_{\mu 4}|^{2}$ in this region arises from the energy dependence of the disappearance signal, which is due to the subleading term $|U_{\mu 4}|^{4}\,e^{-\Gamma_{4}L}$ in $P_{\mu\mu}$ and is maximal for $gm_{4}\sim 10\,\mbox{eV}$. For the appearance channel, instead, the signal strongly depends on $gm_{4}$. For large values of this parameter, corresponding to a large $\Gamma_{4}$, most sterile neutrinos decay before reaching the DUNE near detector and the appearance probability (7) reduces to $P_{\mu e}\simeq|U_{\mu 4}|^{2}$. This explains the upper part of the green long-dashed curve. For smaller values of $gm_{4}$, hence of $\Gamma_{4}$, only a fraction of the sterile neutrinos decay, leading to a suppression of the appearance probability by a factor $(1-e^{-\Gamma_{4}L})$. As a result, the sensitivity to $|U_{\mu 4}|^{2}$ decreases with $gm_{4}$. Finally, for $gm_{4}$ small enough, most sterile neutrinos reach the near detector before decaying, resulting in a strong suppression of appearance events. The sensitivity to $|U_{\mu 4}|^{2}$ is then driven by neutral current and misidentification background events, whose number is suppressed relative to the SM by a factor $\simeq 1-2|U_{\mu 4}|^{2}(1-|U_{\mu 4}|^{2})$ in the low $gm_{4}$ region (as can be seen by taking the limit $e^{-\Gamma_{4}L}\to 1$ in Eqs. (11) and (12)). This explains the lower part of the exclusion curve, as well as the narrow allowed region [^8] This narrow region is also allowed by disappearance data, since both the signal, Eq. (6), and the background, Eq. (12), are suppressed by a factor $\simeq 1-2|U_{\mu 4}|^{2}(1-|U_{\mu 4}|^{2})$ relative to the SM in the limit $e^{-\Gamma_{4}L}\to 1$. close to $|U_{\mu 4}|^{2}=1$. However, as already mentioned, the sensitivity of the DUNE ND-LAr detector in the low $gm_{4}$ region is not competitive with the MINOS upper bound.
IV.4 Dirac versus Majorana neutrinos
@@
IV.5 Comparing the ND-LAr and MicroBooNE/SBN sensitivities to the HDSN parameters
@@
Fig. 8 corresponds to the Majorana case, which was not considered [^10] The authors of Ref. Hostert:2024etd did not consider the Majorana case in order to avoid conflict with the stringent experimental limits on the flux of $\bar{\nu}_{e}$ from the Sun Hostert:2020oui . However, the constraints discussed in Ref. Hostert:2020oui actually apply to the HDSN scenario of Ref. Dentler:2019dhz , in which the electron neutrino contains a $\nu_{4}$ component and the light scalar field $\phi$ can decay to a neutrino-antineutrino pair, leading to $\nu_{4}\to\nu_{i}\phi\to\nu_{i}\nu_{j}\bar{\nu}_{k}$ ($i,j,k=1,2,3$) in the Sun. In the scenario of Ref. deGouvea:2019qre considered in this paper, $\phi$ is extremely light and does not decay to active neutrinos. While decays $\nu_{4}\to\bar{\nu}_{e}\phi$ are possible in the Majorana case, solar neutrinos are produced in a combination of matter Hamiltonian eigenstates that only very weakly mix with $\nu_{4}$, thus evading the constraints on the $\bar{\nu}_{e}$ flux from the Sun. in Ref. Hostert:2024etd . We therefore show the LSND and MiniBooNE allowed regions from Ref. deGouvea:2019qre , as well as the expected sensitivity of the SBN experiment (of which MicroBooNE is one of the three detectors) from the same paper. The same conclusions as in the Dirac case hold. Namely, ND-LAr will be able to probe a much larger region of the HDSN parameter space than SBN, and either to fully exclude this scenario as a solution of the LSND and MiniBooNE anomalies, or to confirm a possible hint of a positive signal in future MicroBooNE, SBND or ICARUS data. Notice that the shape of the MiniBooNE allowed region given by Ref. deGouvea:2019qre and shown in Fig. 8 is significantly different from the one of Ref. Hostert:2024etd , displayed in Fig. 7. This is due to the fact that only appearance data was considered in Ref. deGouvea:2019qre , while both appearance and disappearance data were included in the fit of Ref. Hostert:2024etd . In practice, however, the main difference between the MiniBooNE allowed regions of Figs. 7 and 8 lies in the part of the parameter space that is already excluded by MINOS/MINOS+. The same comment applies to the sensitivity region of the SBN experiment shown in Fig. 8, which does not take into account disappearance data, at variance with the region excluded by MicroBooNE in Fig. 7. One can also notice that, in the large $gm_{4}$ region of the parameter space, where the sensitivity to $|U_{\mu 4}|^{2}$ is driven by appearance data, the MicroBooNE constraint is already as good as the expected sensitivity of the SBN program reported in Ref. deGouvea:2019qre . This suggests that the final sensitivity of SBN will be better than indicated by the pink long-dashed curve of Fig. 8.
@@
Throughout this paper, we assumed that ND-LAr was located on the beam axis, although it will be a movable detector, able to collect data at different angles from the on-axis position. As discussed in Section III.1 and illustrated in Fig. 1, the value of the off-axis angle strongly affects the neutrino flux at the near detector. Namely, the intensity of the flux decreases as the off-axis angle increases, and its peak is shifted towards lower energies. We have checked that this flux reduction weakens the sensitivity of ND-LAr to $|U_{\mu 4}|^{2}$, as could have been anticipated. This is the reason why we did not present any results for off-axis locations of the DUNE ND-LAr detector. On the other hand, future off-axis data may help to constrain the energy dependence of the flux and cross section uncertainties, and make it possible to check how our results are affected by this energy dependence.
V Conclusions

### v2

# Apartado Postal 14-740 07000, Ciudad de Mexico, Mexico

@@ 1 Introduction
The paper is organized as follows. In Section 2, we introduce the heavy decaying sterile neutrino scenario considered in this paper and provide the main formulae used in our analysis. Section 3 briefly presents the specifications of the DUNE liquid argon near detector and describes the simulations performed in our paper. In Section 4, we present our results for the expected sensitivity of ND-LAr to the HDSN parameters, and comment about the possibility to distinguish between Dirac and Majorana neutrinos if this scenario is realized in Nature. Finally, we give our conclusions in Section 5.
2 The heavy decaying sterile neutrino (HDSN) scenario
@@
$P_{\mu\mu}=P_{\bar{\mu}\bar{\mu}}=(1-|U_{\mu 4}|^{2})^{2}+|U_{\mu 4}|^{4}\,e^{-\Gamma_{4}L}\,,\qquad P_{ee}=P_{\bar{e}\bar{e}}=1\ ,\hskip 56.9055pt$ (6)
$P_{\mu e}=P_{\bar{\mu}\bar{e}}=|U_{\mu 4}|^{2}(1-e^{-\Gamma_{4}L})\ .\hskip 199.16928pt$ (7)
@@
$\frac{dP_{\mu e}}{dE_{e}}=\frac{dP_{\bar{\mu}\bar{e}}}{dE_{e}}=|U_{\mu 4}|^{2}(1-e^{-\Gamma^{\rm D}_{4}L})\,\frac{2E_{e}}{E^{2}_{4}}\qquad$ $\text{(Dirac case)}\,$, (9)
$\frac{dP_{\mu e}}{dE_{e}}=\frac{dP_{\bar{\mu}\bar{e}}}{dE_{e}}=\frac{dP_{\mu\bar{e}}}{dE_{e}}=\frac{dP_{\bar{\mu}e}}{dE_{e}}=|U_{\mu 4}|^{2}(1-e^{-\Gamma^{\rm M}_{4}L})\,\frac{E_{e}}{E^{2}_{4}}\qquad$ $\text{(Majorana\ case)}\,.$ (10)
@@
3 Experimental setup and simulation details
@@ 3.2 Simulation details
Table 1: Event rates at the DUNE liquid argon near detector (ND-LAr) expected in the Standard Model, assuming 3.5 years of on-axis data taking in either neutrino or antineutrino mode. The first and second rows correspond to the appearance channel in the neutrino and antineutrino modes, respectively, for which the signal from neutrino oscillations is negligible. The last three columns show the number of intrinsic, misidentification, and neutral current background events, as defined in Section 2. The third and fourth rows correspond to the disappearance channel in the neutrino and antineutrino modes, respectively, and the number of signal events is given separately for the right sign component (e.g. $\nu_{\mu}\rightarrow\nu_{\mu}$ in the neutrino mode) and for the wrong sign component (e.g. $\bar{\nu}_{\mu}\rightarrow\bar{\nu}_{\mu}$ in the neutrino mode) of the neutrino beam.
@@
$\left[N^{\nu_{\mu}\rightarrow\nu_{e}}_{i}\right]^{D,\,CC}$ $=\frac{nT}{4\pi L^{2}}\,\int_{E_{rec}^{i}}^{E_{rec}^{i+1}}\!\!A(E_{rec})\,dE_{rec}\int_{0}^{E_{\nu}^{max}}\!\!\!dE_{\nu_{e}}\,\mathcal{R}(E_{\nu_{e}},\,E_{rec})\,\sigma^{CC}_{\nu_{e}}(E_{\nu_{e}})\hskip 36.98866pt$
@@
$\left[N^{\nu_{\mu}\rightarrow\nu_{\mu}}_{i}\right]^{D,\,CC}$ $=\frac{nT}{4\pi L^{2}}\,\int_{E_{rec}^{i}}^{E_{rec}^{i+1}}\!\!A(E_{rec})\,dE_{rec}\hskip 202.01474pt$
$\times\int_{0}^{E_{\nu}^{max}}\!\!\!dE_{\nu_{\mu}}\,\mathcal{R}(E_{\nu_{\mu}},\,E_{rec})\,\sigma^{CC}_{\nu_{\mu}}(E_{\nu_{\mu}})\,\phi_{\nu_{\mu}}(E_{\nu_{\mu}})\,P^{D}_{\mu\mu}(E_{\nu_{\mu}})\,$, (19)
$\left[N^{\nu_{\mu}}_{i}\right]$ $^{D,\,NC}=\frac{nT}{4\pi L^{2}}\,\int_{E_{rec}^{i}}^{E_{rec}^{i+1}}\!\!A(E_{rec})\,dE_{rec}\hskip 219.08612pt$
$\times\biggl[\int_{0}^{E_{\nu}^{max}}\!\!\!dE_{\nu_{\mu}}\,\mathcal{R}(E_{\nu_{\mu}},E_{rec})\,\sigma^{NC}_{\nu}(E_{\nu_{\mu}})\,\phi_{\nu_{\mu}}(E_{\nu_{\mu}})\,P^{D}_{\mu\mu}(E_{\nu_{\mu}})$
$+\int_{0}^{E_{\nu}^{max}}\!\!\!dE_{\nu_{e}}\,\mathcal{R}(E_{\nu_{e}},E_{rec})\,\sigma^{NC}_{\nu}(E_{\nu_{e}})\int_{E_{\nu_{e}}}^{E_{\nu}^{max}}\!\!\!dE_{\nu_{\mu}}\,\phi_{\nu_{\mu}}(E_{\nu_{\mu}})\,\frac{dP^{D}_{\mu e}}{dE_{\nu_{e}}}\,(E_{\nu_{\mu}},E_{\nu_{e}})\biggr]\,.$ (20)
@@
$\left[N^{\nu_{\mu}\rightarrow\nu_{\mu}}_{i}\right]^{M,\,CC}$ $=\frac{nT}{4\pi L^{2}}\,\int_{E_{rec}^{i}}^{E_{rec}^{i+1}}\!\!A(E_{rec})\,dE_{rec}\hskip 202.01474pt$
$\times\int_{0}^{E_{\nu}^{max}}\!\!\!dE_{\nu_{\mu}}\,\mathcal{R}(E_{\nu_{\mu}},E_{rec})\,\sigma^{CC}_{\nu_{\mu}}(E_{\nu_{\mu}})\,\phi_{\nu_{\mu}}(E_{\nu_{\mu}})\,P^{M}_{\mu\mu}(E_{\nu_{\mu}})\,$, (22)
$\left[N^{\nu_{\mu}}_{i}\right]$ $^{M,\,NC}=\frac{nT}{4\pi L^{2}}\,\int_{E_{rec}^{i}}^{E_{rec}^{i+1}}\!\!A(E_{rec})\,dE_{rec}$
$\times$ $\,\biggl[\int_{0}^{E_{\nu}^{max}}\!\!\!dE_{\nu_{\mu}}\,\mathcal{R}(E_{\nu_{\mu}},E_{rec})\,\sigma^{NC}_{\nu}(E_{\nu_{\mu}})\,\phi_{\nu_{\mu}}(E_{\nu_{\mu}})\,P^{M}_{\mu\mu}(E_{\nu_{\mu}})$
@@
It is straightforward to generalize Eqs. (18) to (23) to the case where the experiment runs in the antineutrino mode. The probabilities appearing in these equations are given by (see Section 2)
$\frac{dP^{D}_{\mu e}}{dE_{e}}=\frac{dP^{D}_{\bar{\mu}\bar{e}}}{dE_{e}}=|U_{\mu 4}|^{2}(1-e^{-\Gamma^{\rm D}_{4}L})\,\frac{2E_{e}}{E^{2}_{4}}\qquad$ $\text{(Dirac case})\,$, (24)
$\frac{dP^{M}_{\mu e}}{dE_{e}}=\frac{dP^{M}_{\bar{\mu}\bar{e}}}{dE_{e}}=\frac{dP^{M}_{\mu\bar{e}}}{dE_{e}}=\frac{dP^{M}_{\bar{\mu}e}}{dE_{e}}=|U_{\mu 4}|^{2}(1-e^{-\Gamma^{\rm M}_{4}L})\,\frac{E_{e}}{E^{2}_{4}}\qquad$ $\text{(Majorana case})\,$, (25)
@@
[…]
$P^{M}_{\mu\mu}=P^{M}_{\bar{\mu}\bar{\mu}}=\left(1-|U_{\mu 4}|^{2}\right)^{2}+|U_{\mu 4}|^{4}e^{-\Gamma^{\rm M}_{4}L}\qquad\text{(Majorana case)}\,.$ (27)
4 ND-LAr sensitivity to the HDSN parameters
@@
4.1 Appearance and disappearance spectra in the Standard Model
@@
4.2 Appearance and disappearance spectra in the HDSN scenario
@@
4.3 Sensitivity of ND-LAr to the HDSN parameters
@@
The different shapes of the appearance and disappearance sensitivity curves in the left plot of Fig. 5 can be understood from the formulae of Section 2. For the disappearance channel, the difference between the predictions of the SM and HDSN scenario for the signal and background scales as $|U_{\mu 4}|^{2}$, as already noticed, with little dependence on $gm_{4}$ (see Eqs. (6) and (12), where the $gm_{4}$ dependence is hidden in the sterile neutrino decay rate $\Gamma_{4}$). In fact, the sensitivity to $|U_{\mu 4}|^{2}$ in the disappearance channel is almost independent of $gm_{4}$, except around $gm_{4}\sim 10\,\mbox{eV}$. The increased sensitivity to $|U_{\mu 4}|^{2}$ in this region arises from the energy dependence of the disappearance signal, which is due to the subleading term $|U_{\mu 4}|^{4}\,e^{-\Gamma_{4}L}$ in $P_{\mu\mu}$ and is maximal for $gm_{4}\sim 10\,\mbox{eV}$. For the appearance channel, instead, the signal strongly depends on $gm_{4}$. For large values of this parameter, corresponding to a large $\Gamma_{4}$, most sterile neutrinos decay before reaching the DUNE near detector and the appearance probability (7) reduces to $P_{\mu e}\simeq|U_{\mu 4}|^{2}$. This explains the upper part of the green long-dashed curve. For smaller values of $gm_{4}$, hence of $\Gamma_{4}$, only a fraction of the sterile neutrinos decay, leading to a suppression of the appearance probability by a factor $(1-e^{-\Gamma_{4}L})$. As a result, the sensitivity to $|U_{\mu 4}|^{2}$ decreases with $gm_{4}$. Finally, for $gm_{4}$ small enough, most sterile neutrinos reach the near detector before decaying, resulting in a strong suppression of appearance events. The sensitivity to $|U_{\mu 4}|^{2}$ is then driven by neutral current and misidentification background events, whose number is suppressed relative to the SM by a factor $\simeq 1-2|U_{\mu 4}|^{2}(1-|U_{\mu 4}|^{2})$ in the low $gm_{4}$ region (as can be seen by taking the limit $e^{-\Gamma_{4}L}\to 1$ in Eqs. (11) and (12)). This explains the lower part of the exclusion curve, as well as the narrow allowed region [^8] This narrow region is also allowed by disappearance data, since both the signal, Eq. (6), and the background, Eq. (12), are suppressed by a factor $\simeq 1-2|U_{\mu 4}|^{2}(1-|U_{\mu 4}|^{2})$ relative to the SM in the limit $e^{-\Gamma_{4}L}\to 1$. close to $|U_{\mu 4}|^{2}=1$. However, as already mentioned, the sensitivity of the DUNE ND-LAr detector in the low $gm_{4}$ region is not competitive with the MINOS upper bound.
4.4 Dirac versus Majorana neutrinos
@@
4.5 Comparing the ND-LAr and MicroBooNE/SBN sensitivities to the HDSN parameters
@@
Fig. 8 corresponds to the Majorana case, which was not considered [^10] The authors of Ref. Hostert:2024etd did not consider the Majorana case in order to avoid conflict with the stringent experimental limits on the flux of $\bar{\nu}_{e}$ from the Sun Hostert:2020oui. However, the constraints discussed in Ref. Hostert:2020oui actually apply to the HDSN scenario of Ref. Dentler:2019dhz, in which the electron neutrino contains a $\nu_{4}$ component and the light scalar field $\phi$ can decay to a neutrino-antineutrino pair, leading to $\nu_{4}\to\nu_{i}\phi\to\nu_{i}\nu_{j}\bar{\nu}_{k}$ ($i,j,k=1,2,3$) in the Sun. In the scenario of Ref. deGouvea:2019qre considered in this paper, $\phi$ is extremely light and does not decay to active neutrinos. While decays $\nu_{4}\to\bar{\nu}_{e}\phi$ are possible in the Majorana case, solar neutrinos are produced in a combination of matter Hamiltonian eigenstates that only very weakly mix with $\nu_{4}$, thus evading the constraints on the $\bar{\nu}_{e}$ flux from the Sun. in Ref. Hostert:2024etd.
@@
We therefore show the LSND and MiniBooNE allowed regions from Ref. deGouvea:2019qre, as well as the expected sensitivity of the SBN experiment (of which MicroBooNE is one of the three detectors) from the same paper. The same conclusions as in the Dirac case hold. Namely, ND-LAr will be able to probe a much larger region of the HDSN parameter space than SBN, and either to fully exclude this scenario as a solution of the LSND and MiniBooNE anomalies, or to confirm a possible hint of a positive signal in future MicroBooNE, SBND or ICARUS data. Notice that the shape of the MiniBooNE allowed region given by Ref. deGouvea:2019qre and shown in Fig. 8 is significantly different from the one of Ref. Hostert:2024etd, displayed in Fig. 7. This is due to the fact that only appearance data was considered in Ref. deGouvea:2019qre, while both appearance and disappearance data were included in the fit of Ref. Hostert:2024etd. In practice, however, the main difference between the MiniBooNE allowed regions of Figs. 7 and 8 lies in the part of the parameter space that is already excluded by MINOS/MINOS+. The same comment applies to the sensitivity region of the SBN experiment shown in Fig. 8, which does not take into account disappearance data, at variance with the region excluded by MicroBooNE in Fig. 7. One can also notice that, in the large $gm_{4}$ region of the parameter space, where the sensitivity to $|U_{\mu 4}|^{2}$ is driven by appearance data, the MicroBooNE constraint is already as good as the expected sensitivity of the SBN program reported in Ref. deGouvea:2019qre. This suggests that the final sensitivity of SBN will be better than indicated by the pink long-dashed curve of Fig. 8.
@@
Throughout this paper, we assumed that ND-LAr was located on the beam axis, although it will be a movable detector, able to collect data at different angles from the on-axis position. As discussed in Section 3.1 and illustrated in Fig. 1, the value of the off-axis angle strongly affects the neutrino flux at the near detector. Namely, the intensity of the flux decreases as the off-axis angle increases, and its peak is shifted towards lower energies. We have checked that this flux reduction weakens the sensitivity of ND-LAr to $|U_{\mu 4}|^{2}$, as could have been anticipated. This is the reason why we did not present any results for off-axis locations of the DUNE ND-LAr detector. On the other hand, future off-axis data may help to constrain the energy dependence of the flux and cross section uncertainties, and make it possible to check how our results are affected by this energy dependence.
5 Conclusions